# Mining experimental data from Materials Science literature with Large Language Models

### v1

@@ A.3.2 Few-shot
Run1 Strict matching 67.97 50.31 57.82 846
Soft matching 69.27 51.27 58.92 846
Run2 Strict matching 68.2 50.48 58.02 846
Soft matching 69.5 51.44 59.13 846
Run3 Strict matching 69.23 50.39 58.33 832
Soft matching 71.18 51.09 59.14 832
@@
Strict matching 58.05 0.25 841
Soft matching 59.06 0.12
@@
Run1 Strict matching 63.97 47.68 54.64 852
Soft matching 66.08 49.26 56.44 852
Run2 Strict matching 66.02 47.59 55.31 824
Soft matching 67.23 48.47 56.33 824
Run3 Strict matching 62.24 45.14 52.33 829
Soft matching 63.93 46.37 53.75 829
@@
Strict matching 54.09 1.56 835
Soft matching 55.50 1.52
Table 24: Performance Metrics for the GPT4 model on RE in materials-properties extraction. In this run, entities are not shuffled when they are added to the prompt. P: Precision, R: Recall, F1: harmonic average of P and R, Supp: Support, number of extracted entities.
@@
Run1 Strict matching 80.57 71.83 75.95 1019
Soft matching 80.57 71.83 75.95 1019
Run2 Strict matching 81.29 72.62 76.71 1021
Soft matching 81.39 72.7 76.8 1021
Run3 Strict matching 80.9 72.27 76.34 1021
Soft matching 80.9 72.27 76.34 1021
@@
Strict matching 76.33 0.38 1020
Soft matching 76.36 0.42
@@
Run1 Strict matching 80.57 70.17 74.5 1010
Soft matching 80.57 70.25 74.59 1010
Run2 Strict matching 81.29 70.87 75.03 1016
Soft matching 81.39 71.13 75.31 1016
Run3 Strict matching 80.9 70.6 75.07 1007
Soft matching 80.9 70.6 75.07 1007
@@
Strict matching 74.86 0.31 1011
Soft matching 74.99 0.36
@@
Run1 Strict matching 83.44 70.08 76.18 960
Soft matching 84.69 71.13 77.32 960
Run2 Strict matching 84.1 70.34 76.61 956
Soft matching 85.46 71.48 77.85 956
Run3 Strict matching 84.01 70.34 76.57 957
Soft matching 85.37 71.48 77.81 957
@@
Strict matching 76.45 0.23 957
Soft matching 77.66 0.29
@@
Run1 Strict matching 82.01 67.02 73.76 934
Soft matching 83.51 68.24 75.11 934
Run2 Strict matching 84.11 69.47 76.09 944
Soft matching 85.49 70.6 77.34 944
Run3 Strict matching 82.31 67.98 74.46 944
Soft matching 84.32 69.64 76.28 944
@@
Strict matching 74.77 1.19 940
Soft matching 76.24 1.11

### v3

@@ A.3.2 Few-shot
Run1 Strict matching 76.38 70.17 73.14 1050
Soft matching 76.95 70.69 73.69 1050
Run2 Strict matching 75.78 69.82 72.68 1053
Soft matching 76.35 70.34 73.22 1053
Run3 Strict matching 75.93 69.82 72.74 1051
Soft matching 76.78 70.6 73.56 1051
@@
Strict matching 72.85 0.25 1051
Soft matching 73.49 0.24
@@
Run1 Strict matching 72.66 66.49 69.44 1046
Soft matching 73.52 67.28 70.26 1046
Run2 Strict matching 72.61 65.18 68.7 1026
Soft matching 74.07 66.49 70.08 1026
Run3 Strict matching 73.16 66.05 69.43 1032
Soft matching 74.42 67.19 70.62 1032
@@
Strict matching 72.85 0.25 1051
Soft matching 73.49 0.24
Table 24: Performance Metrics for the GPT4 model on RE in materials-properties extraction. In this run, entities are not shuffled when added to the prompt. P: Precision, R: Recall, F1: harmonic average of P and R, Supp: Support, number of extracted entities.
@@
Run1 Strict matching 82.41 74.19 78.08 1029
Soft matching 82.41 74.19 78.08 1029
Run2 Strict matching 83.32 73.84 78.29 1013
Soft matching 83.51 74.02 78.48 1013
Run3 Strict matching 83.5 73.93 78.42 1012
Soft matching 83.6 74.02 78.52 1012
@@
Strict matching 78.26 0.17 1018
Soft matching 78.36 0.24
@@
Run1 Strict matching 81.19 72.88 76.81 1026
Soft matching 81.19 72.88 76.81 1026
Run2 Strict matching 81.19 72.88 76.81 1026
Soft matching 81.19 72.88 76.81 1026
Run3 Strict matching 82.77 74.37 78.34 1027
Soft matching 82.86 74.45 78.43 1027
@@
Strict matching 77.46 0.78 1022
Soft matching 77.52 0.80
@@
Run1 Strict matching 86.48 71.65 78.37 947
Soft matching 87.54 72.53 79.33 947
Run2 Strict matching 86.01 71.57 78.13 951
Soft matching 86.01 71.57 78.13 951
Run3 Strict matching 86.67 71.13 78.14 938
Soft matching 87.53 71.83 78.9 938
@@
Strict matching 78.21 0.13 945
Soft matching 79.04 0.25
@@
Run1 Strict matching 86.59 70.08 77.47 925
Soft matching 87.46 70.78 78.24 925
Run2 Strict matching 87.17 70.17 77.75 920
Soft matching 87.93 70.78 78.43 920
Run3 Strict matching 86.15 69.12 76.7 917
Soft matching 87.02 69.82 77.48 917
@@
Strict matching 77.30 0.54 920
Soft matching 78.05 0.50